# On the principle of linearized stability for quasilinear evolution equations in time-weighted spaces

### v1

@@ 1. Introduction
assuming a strict inclusion $\mathrm{dom}(f)\subsetneq\mathrm{dom}(A)$ for the domains of the function $v\mapsto f(v)$ and the quasilinear part $v\mapsto A(v)$. We shall establish herein the principle of linearized stability in phase spaces for the initial values that lie between $\mathrm{dom}(f)$ and $\mathrm{dom}(A)$ enabling greater flexibility in applications. For this purpose we rely on previous results [14, 12] on the well-posedness of (1.1) in time-weighted function spaces.
@@
$O_{\beta}\,\text{ is an open subset of }E_{\beta}\,.$ (1.2c)
@@
Moreover, we define a critical value
$\alpha_{\rm crit}:=\frac{q\xi-1-\gamma}{q-1}\quad\text{ if }q>1\,,\qquad\alpha_{\rm crit}:=-\infty\quad\text{ if }q=1\,,$ (1.2e)
with $\alpha_{\rm crit}<\xi$, and consider initial values $v^{0}\in O_{\alpha}:=O_{\beta}\cap E_{\alpha}$ with
@@
The well-posedness theory for (1.1) developed in [12] in the critical case $\alpha=\alpha_{\rm crit}\in(\beta,\xi)$ relies on the assumption that there are interpolation functors {,}$\{\cdot,\cdot\}_{\alpha/\xi}$ and {,}${\{\cdot,\cdot\}_{\gamma/\eta}}$ of exponents $\alpha/\xi$ and $\gamma/\eta$, respectively, for $\eta\in\{\alpha,\beta,\xi\}$, such that
@@
Assumption (1.3) is automatically satisfied if, for each $\theta\in\{\gamma,\beta,\alpha,\xi\}$, the interpolation functor ${(\cdot,\cdot)_{\theta}}$ is chosen to be always either the complex interpolation functor $[\cdot,\cdot]_{\theta}$, or the continuous interpolation functor ${(\cdot,\cdot)_{\theta,\infty}^{0}}$, or the real interpolation functor $(\cdot,\cdot)_{\theta}=(\cdot,\cdot)_{\theta,p}$ with parameter $p\in[1,\infty]$. This is a consequence of the reiteration theorems for these functors, see, e.g., [4, I.Remarks 2.11.2 (b)].
@@
The special case (1.4) of (1.2d) frequently arises in applications; see, for instance, the examples provided in Section 4.
Let us point out with respect to the parameters in (1.2a) that the parameter $q\geq 1$ measures the growth of the nonlinearity $f$ with respect to the $E_{\xi}$-terms, and its value is generally fixed and usually cannot be adjusted freely. In certain applications, such as considered in Examples 4.2 and 4.3, there may be, however, some flexibility with respect to the choice of the interpolation exponents $\,\beta,\gamma$, and $\xi$. These parameters then define the range of ${\alpha\geq\alpha_{\rm crit}}$ for the phase space $E_{\alpha}$ with critical value $\alpha=\alpha_{\rm crit}$. In certain applications, the critical space $E_{\alpha_{\rm crit}}$ can also be identified due to scaling invariance properties (see Example 4.3).
For simplicity, we only consider herein the situation that $\alpha\in(\beta,\xi)$ since the principle of linearized stability for the case $\alpha>\xi$ is treated in [13]. We emphasize that $E_{\xi}\hookrightarrow E_{\alpha}\hookrightarrow E_{\beta}$ in the setting herein and hence, the semilinear part $f$, being defined on $O_{\xi}$, does not necessarily need to be defined on the phase space $E_{\alpha}$ and requires possibly more regularity than the quasilinear part $A$.
@@
The existence and uniqueness of the maximal strong solution is established in [12, Theorem 1.2] for the critical value $\alpha_{\rm crit}=\alpha\in(\beta,\xi)$. Moreover, in the noncritical case $\alpha_{\rm crit}<\alpha\in(\beta,\xi)$, it is proven in [14, Theorem 1.1] (without assuming (1.3)) that (1.1) has a unique maximal strong solution which satisfies (1.5a) together with
@@
We shall now present the principle of linearized stability in the phase space $E_{\alpha}$ for a certain range of exponents $\alpha$. With regard to applications it proves crucial to state sharp results including certain special cases. To this end, let
@@
with Fréchet derivatives $\partial f(v_{*})\in\mathcal{L}(E_{\xi},E_{0})$ and $(\partial A(v_{*})[\cdot])v_{*}\in\mathcal{L}(E_{\xi},E_{0})$, respectively, such that there are
$\gamma_{*}\in[0,\gamma]\,,\qquad q_{*}>1\,,$ (1.8c)
and $r_{*},\,c_{*}>0$ with
$\|f(w+v_{*})-f(v_{*})-\partial f(v_{*})w\|_{\gamma_{*}}\leq c_{*}\|w\|_{\xi}^{q_{*}}\,,\qquad w\in\widehat{O}_{\xi}\cap{\overline{\mathbb{B}}_{E_{\alpha}}(0,r_{*})}\,,$ (1.8d)
and
$\|A(w+v_{*})v_{*}-A(v_{*})v_{*}-(\partial A(v_{*})[w])v_{*}\|_{\gamma_{*}}\leq c_{*}\|w\|_{\xi}^{q_{*}}\,,\qquad w\in\widehat{O}_{\xi}\cap{\overline{\mathbb{B}}_{E_{\alpha}}(0,r_{*})}\,,$ (1.8e)
where $\widehat{O}_{\xi}:=O_{\xi}-v_{*}$. Moreover, we assume that the linearized operator
@@
$-\omega_{0}:=s(\mathbb{A}):=\sup\left\{\mathrm{Re}\,\lambda\,:\,\lambda\in\sigma(\mathbb{A})\right\}<0\,.$ (1.8f)
We shall prove the following result regarding the asymptotic exponential stability of the equilibrium $v_{*}$.
@@
Assume (1.2), (1.3), and (1.8) and set
$\alpha_{\rm crit}^{*}:=\frac{q_{*}\xi-1-\gamma_{*}}{q_{*}-1}<\xi\,.$
Moreover, assume $\alpha\geq\alpha_{\rm crit}^{*}$ with strict inequality $\alpha>\alpha_{\rm crit}^{*}$ in case that $\gamma_{*}\in(0,\gamma)$. Then, the equilibrium $v_{*}$ is asymptotically exponentially stable in $E_{\alpha}$. More precisely, given any $\omega\in(0,\omega_{0})$, there exist $\varepsilon_{0}>0$ and ${M\geq 1}$ such that, for each $v^{0}\in\overline{\mathbb{B}}_{E_{\alpha}}(v_{*},\varepsilon_{0})$, the solution to (1.1) exists globally in time and
$\|v(t;v^{0})-v_{*}\|_{\alpha}+t^{\xi-\alpha}\|v(t;v^{0})-v_{*}\|_{\xi}\leq Me^{-\omega t}\|v^{0}-v_{*}\|_{\alpha}\,,\qquad t\geq 0\,.$ (1.9)
@@
In the noncritical case $\max\{\alpha_{\rm crit},\,\alpha_{\rm crit}^{*}\}<\alpha\in(\beta,\xi)$ one may actually drop the assumption (1.3) in Theorem 1.3. The claim remains valid (with a similar proof) provided (1.9) is replaced by
$\|v(t;v^{0})-v_{*}\|_{\alpha}+t^{\mu}\|v(t;v^{0})-v_{*}\|_{\xi}\leq Me^{-\omega t}\|v^{0}-v_{*}\|_{\alpha}\,,\qquad t\geq 0\,,$
for some fixed (but arbitrary) $\mu>\xi-\alpha$. Since in many applications (1.3) is automatically satisfied (see Remark 1.1), we chose to present the (slightly) less general but more concise result in Theorem 1.3.
With respect to the assumption (1.8e), note that, if $v_{*}=0$ and $\partial f(v_{*})=0$, then one may take, in the particular context of (1.4), $\gamma_{*}=\gamma$ and ${q_{*}=q}$ in (1.8c)-(1.8e) (and hence $\alpha_{\rm crit}=\alpha_{\rm crit}^{*}$). This observation proves very useful in the critical case, as illustrated in Example 4.3. Moreover, even if $v_{*}$ is non-zero, since ${A(v_{*})v_{*}=-f(v_{*})\in E_{\gamma}}$ by (1.8a), the equilibrium $v_{*}$ (and consequently the linearizations $(\partial A(v_{*})[\cdot])v_{*}$ and  $\partial f(v_{*})$) may exhibit higher regularity properties in applications. As a consequence, one may more broadly expect to have $\gamma_{*}>0$ in (1.8c)-(1.8e).
We also point out that though $(\gamma_{*},q_{*})$ and $(\gamma,q)$ need not be related, in applications it often turns out that $(\gamma_{*},q_{*})=(\gamma,q)$ and therefore $\alpha_{\rm crit}=\alpha_{\rm crit}^{*}$, see the examples in Section 4.
@@
and define again
$\alpha_{\rm crit}:=\frac{q\xi-1-\gamma}{q-1}\quad\text{ if }q>1\,,\qquad\alpha_{\rm crit}:=-\infty\quad\text{ if }q=1\,,$ (1.11c)
noticing that $\alpha_{\rm crit}<\xi$. We assume that
@@
and let $O_{\alpha}$ be an arbitrary open subset of $E_{\alpha}$. The semilinearity $f:O_{\xi}:=O_{\alpha}\cap E_{\xi}\to E_{\gamma}$ is again assumed to be locally Lipschitz continuous in the sense that for each ${R>0}$ there is a constant $c(R)>0$ such that
@@
The well-posedness of (1.1) is established in [12] for the critical case $\alpha=\alpha_{\rm crit}\in(\gamma,\xi)$ under the assumption that, if $\xi<1$, there exists an interpolation functor {,}$\{\cdot,\cdot\}_{\alpha/\xi}$ of exponent $\alpha/\xi$, and, if ${\gamma>0}$, that for ${\eta\in\{\alpha,\xi\}\setminus\{1\}}$ there are interpolation functors {,}$\{\cdot,\cdot\}_{\gamma/\eta}$ of exponent $\gamma/\eta$ such that
@@
Under these assumptions the Cauchy problem (1.10) is locally well-posed in $O_{\alpha}$. Similarly to the quasilinear problem (1.1), we also assume in the noncritical case the interpolation property (1.12). This leads to a unified and slightly refined well-posedness result for (1.10), compared to [12, Theorem 1.3].
@@
$\underset{t\nearrow t^{+}(v^{0})}{\limsup}\|f(v(t;v^{0}))\|_{0}=\infty\qquad\text{ or }\qquad\lim_{t\nearrow t^{+}(v^{0})}\mathrm{dist}_{E_{\alpha}}\big{(}v(t;v^{0}),\partial O_{\alpha}\big{)}=0$ (1.14)
@@
This result is in [12, Theorem 1.3] in the critical case $\alpha=\alpha_{\rm crit}\in(\gamma,\xi)$. For $\alpha_{\rm crit}<\alpha\in[\gamma,\xi)$ the claim follows by arguing along the lines of the proof of [12, Theorem 1.3] (we omit therefore the details). For an alternative proof in the case $\alpha_{\rm crit}<\alpha\in[\gamma,\xi)$ under the restriction $O_{\alpha}=E_{\alpha}$ (but with a stronger blow-up criterion than in (1.14)) we refer to [14, Theorem 1.2]. ∎
@@
For the Fréchet derivative $\partial f(v_{*})\in\mathcal{L}(E_{\xi},E_{0})$ we assume that there exist
$\gamma_{*}\in[0,\gamma]\,,\qquad q_{*}>1\,,$ (1.15c)
and constants $r_{*},\,c_{*}>0$ such that
$\|f(w+v_{*})-f(v_{*})-\partial f(v_{*})w\|_{\gamma_{*}}\leq c_{*}\|w\|_{\xi}^{q_{*}}\,,\qquad w\in\widehat{O}_{\xi}\cap{\overline{\mathbb{B}}_{E_{\alpha}}(0,r_{*})}\,,$ (1.15d)
where again $\widehat{O}_{\xi}:=O_{\xi}-v_{*}$. Finally, we assume that the linearized operator
@@
has a negative spectral bound, that is
$-\omega_{0}:=s(\mathbb{A}):=\sup\left\{\mathrm{Re}\,\lambda\,:\,\lambda\in\sigma(\mathbb{A})\right\}<0\,.$ (1.15e)
We can now state the exponential stability of the equilibrium solution $v_{*}$ to the semilinear problem (1.10):
@@
Assume (1.11), (1.12), and (1.15) and set
$\alpha_{\rm crit}^{*}:=\frac{q_{*}\xi-1-\gamma_{*}}{q_{*}-1}<\xi\,.$
Moreover, assume $\alpha\geq\alpha_{\rm crit}^{*}$ with strict inequality $\alpha>\alpha_{\rm crit}^{*}$ in case that $\gamma_{*}\in(0,\gamma)$ or $\alpha=\gamma$. Then, the equilibrium $v_{*}$ is asymptotically exponentially stable in $E_{\alpha}$. More precisely, given any ${\omega\in(0,\omega_{0})}$, there are $\varepsilon_{0}>0$ and ${M\geq 1}$ such that, for each $v^{0}\in\overline{\mathbb{B}}_{E_{\alpha}}(v_{*},\varepsilon_{0})$, the solution to (1.10) exists globally in time and
$\|v(t;v^{0})-v_{*}\|_{\alpha}+t^{\xi-\alpha}\|v(t;v^{0})-v_{*}\|_{\xi}\leq Me^{-\omega t}\|v^{0}-v_{*}\|_{\alpha}\,,\qquad t\geq 0\,.$ (1.16)
Remark 1.7.
As for the quasilinear problem (1.1), in the noncritical case $\max\{\alpha_{\rm crit},\,\alpha_{\rm crit}^{*}\}<\alpha\in[\gamma,\xi)$, the assumption (1.12) in Theorem 1.6 can be omitted, and the claim remains valid (with a similar proof) provided that (1.16) is replaced by
$\|v(t;v^{0})-v_{*}\|_{\alpha}+t^{\mu}\|v(t;v^{0})-v_{*}\|_{\xi}\leq Me^{-\omega t}\|v^{0}-v_{*}\|_{\alpha}\,,\qquad t\geq 0\,,$
@@
As in the quasilinear case, in applications it often has $(\gamma_{*},q_{*})=(\gamma,q)$ and $\alpha_{\rm crit}=\alpha_{\rm crit}^{*}$, see the examples in Section 4.
@@
The outline of the paper is as follows. Section 2 is dedicated to the quasilinear problem (1.1). We provide a detailed proof of Theorem 1.3. In Section 3 we sketch the proof of Theorem 1.6 for the semilinear problem (1.10). In Section 4 we present examples that illustrate our previous findings both for the critical case $\alpha=\alpha_{\rm crit}$ and the non-critical case $\alpha>\alpha_{\rm crit}$, respectively, including quasilinear problems with quadratic semilinearities, a parabolic-parabolic chemotaxis system, and a quasilinear evolution equation in critical spaces with scaling invariance.
@@
Before establishing the proofs of Theorem 1.3 and Theorem 1.8 we recall the following consequence of (the proof of) Theorem 1.2 that relies on the semiflow property of the solution map associated with problem (1.1).
Corollary 2.1.
Assume (1.2), (1.3), and (1.8a). Then, given ${t_{*}>1}$, there are a neighborhood $V_{\alpha}$ of $v_{*}$ in $O_{\alpha}$ and $k_{0}\geq 1$ such that $t_{*}<t^{+}(v^{0})$ and
$\|v(t;v^{0})-v_{*}\|_{\alpha}\leq k_{0}\|v^{0}-v_{*}\|_{\alpha}\,,\quad 0\leq t\leq t_{*}\,,$ (2.1)
for each $v^{0}\in V_{\alpha}$. Moreover, there exist $t_{0}\in(0,1)$ and $k_{1}>0$ such that
$\|v(t;v^{0})-v(s;v^{0})\|_{\beta}\leq k_{1}(t-s)^{\alpha-\beta},\quad 0\leq s\leq t\leq t_{0}\,,\quad v^{0}\in V_{\alpha}\,.$ (2.2)
Proof.
Property (2.1) is a consequence of the continuous dependence in $E_{\alpha}$ stated in [14, Theorem 1.1 (iii)] for ${\alpha_{\rm crit}<\alpha\in(\beta,\xi)}$, respectively in [12, Theorem 1.2 (iii)] in the critical case $\alpha_{\rm crit}=\alpha\in(\beta,\xi)$ (see the proof thereof). The second statement is shown in the proof of [14, Proposition 2.1] (see (2.13) therein) for $\alpha_{\rm crit}<\alpha\in(\beta,\xi)$, respectively in the proof of[12, Proposition 2.1] (see (2.19)-(2.22) therein) in the case ${\alpha_{\rm crit}=\alpha\in(\beta,\xi)}$. ∎
The proof of Theorem 1.3 is performed in the framework of time-weighted spaces. Recall that, given a Banach space $E$, $\mu\in\mathbb{R}$, and $T>0$, the space

### v2

@@ 1. Introduction
(and the corresponding semilinear counterparts) assuming a strict inclusion $\mathrm{dom}(f)\subsetneq\mathrm{dom}(A)$ for the domains of the function $v\mapsto f(v)$ and the quasilinear part $v\mapsto A(v)$. We shall establish herein the principle of linearized stability in phase spaces for the initial values that lie between $\mathrm{dom}(f)$ and $\mathrm{dom}(A)$ enabling greater flexibility in applications. For this purpose we rely on previous results [14, 12] on the well-posedness of (1.1) in time-weighted function spaces and extend the principle of linearized stability [13] (for the case $\mathrm{dom}(f)=\mathrm{dom}(A)$) to the present setting.
@@
$\emptyset\not=O_{\beta}\,\text{ is an open subset of }E_{\beta}\,.$ (1.2c)
@@
Furthermore, we define a critical value $\alpha_{\rm crit}<\xi$ by setting
[…]
and consider initial values $v^{0}\in O_{\alpha}:=O_{\beta}\cap E_{\alpha}$ with
@@
The well-posedness theory for (1.1) developed in [12] for the critical case $\alpha=\alpha_{\rm crit}\in(\beta,\xi)$ relies on the assumption that there is an interpolation functor $\{\cdot,\cdot\}_{\alpha/\xi}$ of exponent $\alpha/\xi$ and, for $\eta\in\{\alpha,\beta,\xi\}$, there are interpolation functors $\{\cdot,\cdot\}_{\gamma/\eta}$} of exponents $\gamma/\eta$ such that
@@
Assumption (1.3) is automatically satisfied if, for each $\theta\in\{\gamma,\beta,\alpha,\xi\}$, the interpolation functor $(\cdot,\cdot)_{\theta}$} is chosen to be always either the complex interpolation functor $\cdot,\cdot]_{\theta}$, or the continuous interpolation functor $(\cdot,\cdot)_{\theta,\infty}^{0}$}, or the real interpolation functor $(\cdot,\cdot)_{\theta}=(\cdot,\cdot)_{\theta,p}$ with parameter $p\in[1,\infty]$. This is a consequence of the reiteration theorems for these functors, see e.g. [4, I.Remarks 2.11.2 (b)].
@@
The special case (1.4) of (1.2d) frequently arises in applications; see, for instance, the examples provided in Section 4 and [14, Lemma 4.1].
Let us point out with respect to the parameters in (1.2a) that the parameter $q\geq 1$ measures the growth of the nonlinearity $f$ with respect to the $E_{\xi}$-terms, and its value is usually fixed and cannot be adjusted freely. In certain applications, such as considered in Examples 4.2 and 4.3 below, there may be, however, some flexibility with respect to the choice of the interpolation exponents $\beta,\gamma$, and $\xi$. These parameters then define the range of $\alpha\geq\alpha_{\rm crit}$} for the phase space $E_{\alpha}$ with critical value $\alpha=\alpha_{\rm crit}$. In certain applications, the critical space $E_{\alpha_{\rm crit}}$ can also be identified due to scaling invariance properties (see Example 4.3).
For simplicity, we only consider herein the situation that $\alpha<\xi$, as the case $\alpha>\xi$ is treated in [13] [^1] The case $\alpha=\xi$ could, in principle, be incorporated into the present analysis, but requires a slightly different functional analytic setting since certain statements (for instance (1.5b) below) no longer hold. . We emphasize that thus $E_{\xi}\hookrightarrow E_{\alpha}\hookrightarrow E_{\beta}$ in the setting herein and hence, the semilinear part $f$, being defined on $O_{\xi}$, need not be defined on the phase space $E_{\alpha}$ and may require higher regularity than the quasilinear part $A$. Compared to the classical setting [3, 7, 13], where both $A$ and $f$ are assumed to be locally Lipschitz continuous on $O_{\xi}$, and well-posedness and stability of equilibria are established for initial data $v^{0}\in O_{\alpha}$ with $\alpha>\xi$, we shall use condition (1.2d) (or (1.11e) in the semilinear case) to establish similar results for initial data $v^{0}\in O_{\alpha}$ with $\alpha<\xi$, additionally assuming (1.2e)–(1.2f) (or (1.11b)–(1.11d) for the semilinear case). In fact, condition (1.2d) (or (1.11e)) implies precise quantitative estimates on the underlying semigroups enabling the use of time-weighted spaces in the analysis of the evolution problem, which effectively capture the (singular) behavior of solutions at $t=0$ in $E_{\xi}$.
@@
The existence and uniqueness of a maximal strong solution is established in [12, Theorem 1.2] for the critical value $\alpha_{\rm crit}=\alpha\in(\beta,\xi)$. Moreover, for the noncritical case $\alpha_{\rm crit}<\alpha\in(\beta,\xi)$, it is proven in [14, Theorem 1.1] (without assuming (1.3)) that (1.1) has a unique maximal strong solution which satisfies (1.5a) together with
@@
We shall now present the principle of linearized stability in the phase space $E_{\alpha}$ for the range of exponents $\alpha$ specified in (1.2f). To this end, let
@@
with Fréchet derivatives $\partial f(v_{*})\in\mathcal{L}(E_{\xi},E_{0})$ and $(\partial A(v_{*})[\cdot])v_{*}\in\mathcal{L}(E_{\xi},E_{0})$, respectively. Moreover, we assume that the linearized operator
@@
$\omega_{0}:=s(\mathbb{A}):=\sup\left\{\mathrm{Re}\,\lambda\,:\,\lambda\in\sigma(\mathbb{A})\right\}<0\,.$ (1.8c)
Our assumptions will allow us to rely on the previously established principle of linearized stability from [13]. In fact, we shall prove the following result regarding the asymptotic exponential stability of the equilibrium $v_{*}$ and thereby extend the corresponding result of [13] to the present framework.
@@
Assume (1.2), (1.3), and (1.8). Then, the equilibrium $v_{*}$ is asymptotically exponentially stable in $E_{\alpha}$. More precisely, given any $\omega\in(0,\omega_{0})$, there exist $\varepsilon_{0}>0$ and $M\geq 1$} such that, for each $v^{0}\in\overline{\mathbb{B}}_{E_{\alpha}}(v_{*},\varepsilon_{0})$, the solution to (1.1) exists globally in time and
$\|v(t;v^{0})-v_{*}\|_{\alpha}+t^{\xi-\alpha}\|v(t;v^{0})-v_{*}\|_{\xi}\leq Me^{-\omega t}\|v^{0}-v_{*}\|_{\alpha}\,,\qquad t>0\,.$ (1.9)
The interpolation assumptions (1.3) are not required in any case:
@@
In the noncritical case $\alpha_{\rm crit}<\alpha\in(\beta,\xi)$ one may actually drop the assumptions (1.3) in Theorem 1.3. The claim remains valid (with a similar proof) provided that the estimate (1.9) is replaced by
$\|v(t;v^{0})-v_{*}\|_{\alpha}+t^{\mu}\|v(t;v^{0})-v_{*}\|_{\xi}\leq Me^{-\omega t}\|v^{0}-v_{*}\|_{\alpha}\,,\qquad t>0\,$,
for some fixed (but arbitrary) $\mu>\xi-\alpha$. Since in many applications (1.3) are automatically satisfied (see Remark 1.1), we chose to present the (slightly) less general, but more concise, result in Theorem 1.3.
@@
and define again $\alpha_{\rm crit}<\xi$ by
[…]
We assume that
@@
and let $O_{\alpha}$ be an arbitrary open subset of $E_{\alpha}$. The semilinearity $f:O_{\xi}:=O_{\alpha}\cap E_{\xi}\to E_{\gamma}$ is again assumed to be locally Lipschitz continuous in the sense that for each R>0{R>0} there is a constant $c(R)>0$ such that
@@
The well-posedness of (1.1) is established in [12] for the critical case $\alpha=\alpha_{\rm crit}\in(\gamma,\xi)$ under the assumption that, when $\xi<1$, there exists an interpolation functor $\{\cdot,\cdot\}_{\alpha/\xi}$ of exponent $\alpha/\xi$, and, when $\gamma>0$}, that for $\eta\in\{\alpha,\xi\}\setminus\{1\}$} there are interpolation functors $\{\cdot,\cdot\}_{\gamma/\eta}$ of exponent $\gamma/\eta$ such that
@@
Under these assumptions the Cauchy problem (1.10) is locally well-posed in $O_{\alpha}$. Similar to the quasilinear problem (1.1), we also assume in the noncritical case the interpolation property (1.12). This leads to a unified and slightly refined well-posedness result for (1.10), compared to [12, Theorem 1.3].
@@
$\underset{t\nearrow t^{+}(v^{0})}{\limsup}\|f(v(t;v^{0}))\|_{0}=\infty\qquad\text{ or }\qquad\lim_{t\nearrow t^{+}(v^{0})}\mathrm{dist}_{E_{\alpha}}\big(v(t;v^{0}),\partial O_{\alpha}\big)=0\,.$ (1.14)
@@
This result is proven in [12, Theorem 1.3] for the critical case $\alpha=\alpha_{\rm crit}\in(\gamma,\xi)$. For $\alpha_{\rm crit}<\alpha\in[\gamma,\xi)$ the claim follows by arguing along the lines of the proof of [12, Theorem 1.3] (we omit therefore the details). For an alternative proof in the case $\alpha_{\rm crit}<\alpha\in[\gamma,\xi)$ under the restriction $O_{\alpha}=E_{\alpha}$ (but with a stronger blow-up criterion than in (1.14)) we refer to [14, Theorem 1.2]. ∎
@@
with Fréchet derivative $\partial f(v_{*})\in\mathcal{L}(E_{\xi},E_{0})$. For the linearized operator
@@
we assume a negative spectral bound, that is,
$\omega_{0}:=s(\mathbb{A}):=\sup\left\{\mathrm{Re}\,\lambda\,:\,\lambda\in\sigma(\mathbb{A})\right\}<0\,.$ (1.15c)
If $0<\gamma<\xi<1$ in (1.11), the principle of linearized stability from [13] can be applied (similarly as in the quasilinear case) to establish the exponential stability of the equilibrium solution $v_{*}$ to the semilinear evolution problem (1.10). In contrast, when $\gamma=0$ or $\xi=1$ in (1.11), this approach is no longer applicable. Nonetheless, exponential stability of the equilibrium $v_{*}$ can still be obtained under additional, yet mild, structural assumptions. More precisely, we assume that there exist
$\gamma_{*}\in[0,\gamma]\,,\qquad q_{*}>1\,$, (1.16a)
together with constants $r_{*},\,c_{*}>0$ such that
$\|f(w+v_{*})-f(v_{*})-\partial f(v_{*})w\|_{\gamma_{*}}\leq c_{*}\|w\|_{\xi}^{q_{*}}\,,\qquad w\in\widehat{O}_{\xi}\cap{\overline{\mathbb{B}}_{E_{\alpha}}(0,r_{*})}\,$, (1.16b)
where $\widehat{O}_{\xi}:=O_{\xi}-v_{*}$. Concerning the choice of $q_{*}$ and $\gamma_{*}$ see also Remark 1.7 below.
@@
Assume (1.12), (1.15), and either
  • (a)
    (1.11) with $0<\gamma<\xi<1$;
or
  • (b)
    (1.11) with $\gamma=0$ or $\xi=1$, and (1.16). In this case set
    $\alpha_{\rm crit}^{*}:=\frac{q_{*}\xi-1-\gamma_{*}}{q_{*}-1}<\xi$
    and assume additionally that $\alpha\geq\alpha_{\rm crit}^{*}$ with strict inequality $\alpha>\alpha_{\rm crit}^{*}$ when $\gamma_{*}\in(0,\gamma)$ or $\alpha=\gamma$.
Then, the equilibrium solution $v_{*}$ to (1.10) is asymptotically exponentially stable in $E_{\alpha}$. That is, given any $\omega\in(0,\omega_{0})$}, there are $\varepsilon_{0}>0$ and $M\geq 1$} such that, for each $v^{0}\in\overline{\mathbb{B}}_{E_{\alpha}}(v_{*},\varepsilon_{0})$, the solution to (1.10) exists globally in time and
$\|v(t;v^{0})-v_{*}\|_{\alpha}+t^{\xi-\alpha}\|v(t;v^{0})-v_{*}\|_{\xi}\leq Me^{-\omega t}\|v^{0}-v_{*}\|_{\alpha}\,,\qquad t>0\,.$ (1.17)
Remarks 1.7.
(i) In applications, it often holds that $(\gamma_{*},q_{*})=(\gamma,q)$ and hence $\alpha_{\rm crit}=\alpha_{\rm crit}^{*}$, see e.g. Example 4.1 below and [12, Example 4.1].
(ii) In the noncritical case $\alpha_{\rm crit}<\alpha=\gamma$ (assuming (a)) or $\max\{\alpha_{\rm crit},\,\alpha_{\rm crit}^{*}\}<\alpha\in[\gamma,\xi)$ (assuming (b)), the condition (1.12) in Theorem 1.6 can be omitted, and the claim remains valid (with a similar proof) provided that (1.17) is replaced by
$\|v(t;v^{0})-v_{*}\|_{\alpha}+t^{\mu}\|v(t;v^{0})-v_{*}\|_{\xi}\leq Me^{-\omega t}\|v^{0}-v_{*}\|_{\alpha}\,,\qquad t>0\,$,
@@
The outline of the paper is as follows. Section 2 is dedicated to the quasilinear problem (1.1), where we provide a proof of Theorem 1.3 and establish the instability result stated in Theorem 1.8. In Section 3 we prove Theorem 1.6 for the semilinear problem (1.10). In Section 4 we present examples that illustrate our previous findings both for the critical case $\alpha=\alpha_{\rm crit}$ and the noncritical case $\alpha>\alpha_{\rm crit}$, respectively, including quasilinear problems with quadratic semilinearities, a parabolic-parabolic chemotaxis system, and a quasilinear evolution equation in critical spaces with scaling invariance.
@@
We shall use the framework of time-weighted spaces, which is well suited to capturing the behavior of solutions to (1.1) (and (1.10)) near 0. Recall that, given a Banach space $E$, $\mu\in\mathbb{R}$, and T>0{T>0}, the space